OpenStudy (anonymous):

What are the exact solutions of x^2 − x − 4 = 0?

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (paxpolaris):

\[x={-b \pm \sqrt{b^2-4ac}\over2a}\]

OpenStudy (paxpolaris):

\[{-(-1) \pm \sqrt{(-1)^2-4\cdot1\cdot(-4)}\over 2\cdot1}={1\pm \sqrt{1+16}\over2}\]

OpenStudy (paxpolaris):

\[={1 \pm \sqrt{17}\over2}\]

OpenStudy (anonymous):

Thanks